OpenStudy (pokemon23):

Find two consecutive integers whose sum is: 91 Find two consecutive integers whose sum is: 91 @Mathematics

OpenStudy (anonymous):

guess and check for this one

OpenStudy (mertsj):

The integers are x and x+1 Their sum is x + x+1 which equals 91 2x+1 = 91. 2x = 90 x = 45 x+1 =46

OpenStudy (anonymous):

x + x + 1 = 91 2x + 1 = 91 2x = 90 x = 45 x + 1 = 46

OpenStudy (anonymous):

someone give me a medal

OpenStudy (anonymous):

44 + 45 = 89 50 + 51 = 101 47 + 48 = 95 45 + 46 = 91 got is

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw

OpenStudy (pokemon23):

ty

OpenStudy (anonymous):

yw